A gas occupies 19.3 L at STP. What volume will it occupy at 735 torr and 57°C?


1
Expert's answer
2022-05-13T03:20:05-0400

Given:

STP (T = 273.15 K; P = 1 atm)

P1 = 1 atm

V1 = 19.3 L

T1 = 273.15 K

P2 = 735 torr × (1 atm / 760 torr) = 0.9671 atm

V2 = unknown

T2 = 57°C + 273.15 = 330.15 K


Solution:

Combined Gas Law can be used:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (1 atm × 19.3 L × 330.15 K) / (0.9671 atm × 273.15 K) = 24.1 L

V2 = 24.1 L


Answer: 24.1 liters